HUBBLE SPACE TELESCOPE DAILY REPORT #1594 PERIOD COVERED: 0000Z (UTC) 03/19/96 - 0000Z (UTC) 03/20/96 Daily Status Report as of 080/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were = reported. 1.2 Completed Two Sets of WF/PC-2 6352 (Dwarf Eliptical Galaxy = Snapshots) The WF/PC-2 was used to survey VCC1876, from a sample of 30 dwarf = eliptical galaxies. The observations were completed as planned, and no = anomalies were reported. 1.3 Completed FGS 6051 (Dynamical Masses For The Stars in the Pre-Main = Sequence Spectroscopic Class) The FGS were used to make astrometric measurements of the location = of star STAR-04560 (aka, GSC2387+535) to measure its relative orbit in a = pre-main sequence binary system. The observations were executed as = scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.5 Completed WF/PC-2 6365 (A Snapshot Survey For Companions and = Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to make snapshot observations of Wolf-Rayet = star WR153in order to look for companion stars and clusters. Photometry = was performed on the Wolf-Rayet star and associated stars. The observatio= ns were completed as planned, and no anomalies were reported. 1.6 Completed GHRS/2 6850 (GHRS G270M Y Deflection Test) The High Resolution Spectrograph (Side Two) was used to make an = internal wavelength calibration observation using grading G270M and lamp = SC2 as part of a Y-deflection test. The observations were completed as = Scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 6003 (The Evolution of Elliptical Galaxies in = distant Clusters) The WF/PC-2 was used to make observations of elliptical galaxies = in distant galactic cluster Gal-Clus-1 to study the evolution of the = structure of these elliptical galaxies. The observations were completed = as planned, and no problems were reported. 1.8 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flat) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.9 Completed FOS/RD 6019 (Supersonic Velocity Dispersion in Dwarf = Nova Accretion Disks) The Faint Object Spectrograph (Red Side) was used to make spectrogr= aphic observations of the dwarf nova OY-Car in order to study its = supersonic velocity dispersion. The intent is to nail down the physical = support mechanism for its accretion disk of this short period eclipsing = dwarf nova. As documented in HSTAR 5537, near the beginning of this = nine-hour proposal, at 079/2234Z, there was an FOS mode 2 target = acquisition failure. Otherwise, the observations were executed as = scheduled, and any affects of the acquisition failure were not reported. 1.10 Completed WF/PC-2 6250 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor) The WF/PC-2 was used to make a series on internal calibration = observations. Internal flats, biases, and K-Spot calibration images were = taken. These observations are designed to monitor the health of the = cameras. The observations were executed as scheduled, and no anomalies = were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: There were no operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: ESTR re-conditioning is scheduled to occur both today and = tomorrow. /DMH